You purchased 3,600 shares in the New Pacific Growth Fund on January 2, 2016, at an offering price of $29.30 per share. The front-end load for this fund is 6 percent, and the back-end load for redemptions within one year is 1 percent. The underlying assets in this mutual fund appreciate (including reinvested dividends) by 6 percent during 2016, and you sell back your shares at the end of the year. If the operating expense ratio for the New Pacific Growth Fund is 1.79 percent, what is your total return from this investment? (Assume that the operating expense is netted against the fund’s return.) (A negative value should be indicated by a minus sign. Do not round intermediate calculations. Enter your answer as a percent rounded to 2 decimal places.)
Answer #1

Initial NAV = Offer Price * (1 - Front-end load %) = $29.3(1 - 0.06) = $27.542

Final NAV = Initial NAV * [1 + (Growth% - Operating Expense Ratio)]

= $27.542[1 + (0.06 - 0.0179)] = $28.70

Sale proceeds per share = Final NAV * (1 - Back-end load%) = $28.70(1 - 0.01) = $28.41

Total return = (Sale proceeds per share - Offer Price) / Offer Price

= ($28.41-29.3) / $29.3 = -$0.8855 / $29.3 = -0.0302, or -3.02%